OpenStudy (anonymous):

Determine all values of x, (if any), at which the graph of the function has a horizontal tangent y(x)=x^(4)-256x+2

OpenStudy (anonymous):

\[y'=4x^3-256\] \[4x^3-256=0\] \[4x^3=256\] \[x^3=64\] \[x=4\]

OpenStudy (anonymous):

Thank youuu.